OpenStudy (anonymous):

Evaluate the expression when x = 32 and y = 2. x -- 4y a 1 -16 b 16 __ 21 c 2 d 4

Mehek (mehek14):

which one did you get?

Mehek (mehek14):

\(\dfrac{32}{4*2}=\)

OpenStudy (anonymous):

i got 1 over 16?

Mehek (mehek14):

how?

Mehek (mehek14):

what's 4*2 = ?

OpenStudy (anonymous):

i divided eight by thirty two

Mehek (mehek14):

it's 32/8 not 8/32

Mehek (mehek14):

how many times does 8 go into 32?

OpenStudy (anonymous):

OHHH! then it's four

Mehek (mehek14):

yea :P

OpenStudy (anonymous):

thank you :)

Mehek (mehek14):

np ^_^

OpenStudy (anonymous):

:CLOSED: